OpenStudy (anonymous):

Consider the functions f(x) = 3x − 2 and g(x) = x2 − 1. What is the value of (f − g)(−2)? −8 −5 −11 3

OpenStudy (mathstudent55):

Once again, first calculate: \(f(-2) = 3(-2) - 2 = ~?\) \(g(-2) = (-2)^2 - 1 = ~?\) Then subtract the second one from the first one.

OpenStudy (anonymous):

so -11

OpenStudy (mathstudent55):

Yes. -8 - 3 = -11 Correct!

OpenStudy (anonymous):

tY

OpenStudy (anonymous):

One more please

OpenStudy (anonymous):

Choose the best description of how to calculate f[g(3)] if f(x) = 3x − 2 and g(x) = x2 − 1. Calculate f(3) and g(3), then multiply them together. Calculate f(3) and g(3), then write the results as an ordered pair. Calculate f(3) and g(3), then subtract them. Calculate g(3), plug the result into f(x) and simplify

OpenStudy (mathstudent55):

Notice that here, g(3) is inside f[g(3)] First you need to calculate g(3) and get a number. Let's call it k. Then you calculate f(k). In other words, whatever number you get from g(3) you plug in as x into f(x).